OpenStudy (darkbluechocobo):

Find the missing value. y is directly proportional to x and inversely proportional to the square of z. y = 20 when x = 50 and z = 5. Find y when x = 3 and z = 6.

OpenStudy (darkbluechocobo):

x=y/z^2 50=20/25

OpenStudy (anonymous):

You're very close. Proportional means that there's a constant, k, there. \[ y = k \frac{x}{z^2} \]

OpenStudy (darkbluechocobo):

buht i thought that y would be k?

OpenStudy (anonymous):

This can't be the case because k is a constant and y is a variable.

OpenStudy (darkbluechocobo):

oh

OpenStudy (anonymous):

You have to use the given information (y=20, x=50, z=5) to solve for k. THEN find y when x=3 and z=6. Does that make sense?

OpenStudy (darkbluechocobo):

yes

OpenStudy (darkbluechocobo):

so then 20=50/2500

OpenStudy (darkbluechocobo):

wait nvm i got that wrong i thought it said x^2 oops

OpenStudy (anonymous):

Also, I noticed you switched x and y in your original equation. A tip to avoid this next time: when it says "is," that's when you know to add an equals sign! So "y is direction proportional to x" would give you y=kx :)

OpenStudy (darkbluechocobo):

oooooooooooo i see :3 thank you

OpenStudy (darkbluechocobo):

20=50/25

OpenStudy (darkbluechocobo):

50/25 is 2 so 20=2 10=k

OpenStudy (anonymous):

You're forgetting the constant k again! lol \[ 20 = k \frac{50}{25} \] What is k?

OpenStudy (darkbluechocobo):

10 is k?

OpenStudy (anonymous):

Yup. So your new equation relating x, y, and z is \[ y = 10 \frac{x}{z^2} \]

OpenStudy (darkbluechocobo):

y=10*3/36

OpenStudy (anonymous):

correct :)

OpenStudy (darkbluechocobo):

5/6 is the answer?

OpenStudy (anonymous):

Mhmm, because that's what y equals when x=3 and z=6.

OpenStudy (darkbluechocobo):

hue yay could I do one more?

OpenStudy (anonymous):

Sure!

OpenStudy (darkbluechocobo):

Find the missing value. x varies inversely as y . x = 5 when y = 50. Find x when y = 20.

OpenStudy (anonymous):

How do you think you should set that up?

OpenStudy (darkbluechocobo):

well it cant be like last time so x=k/y?

OpenStudy (darkbluechocobo):

well we have to have squared thing T-T i dont know

OpenStudy (anonymous):

What squared thing? If the statement is "x varies inversely with y," you did it correctly

OpenStudy (darkbluechocobo):

i mean what we did with the last one z^2

OpenStudy (darkbluechocobo):

i thought there is always a squared variable so iset it up correctly then?

OpenStudy (anonymous):

There's only a squared variable if it says so in the statement, as in "x varies with the square of y," but there's no square in this problem, so you did it correctly.

OpenStudy (darkbluechocobo):

hokay

OpenStudy (darkbluechocobo):

5=k/50

OpenStudy (anonymous):

Good

OpenStudy (darkbluechocobo):

(5*50) x= k/20

OpenStudy (anonymous):

Not quite. Remember you want to use the first set of information (x=5, y=50) to solve for k! Then plug this k value into x=k/y again

OpenStudy (darkbluechocobo):

im confused can you show me how you plug them in is it like 5=k/50*20

OpenStudy (anonymous):

Before you plug in for y=20 and try to solve for x, you need to find k. This is because if we went straight to plugging in, you'd have \[ x = k/20 \] and you can't solve that! Therefore, you're using the equation you just came up with, \[5=k/50\] to solve for k. What is the value of k?

OpenStudy (darkbluechocobo):

k=5*50?

OpenStudy (anonymous):

Yup, which is 250. Now that you have \[k=250\] you can plug k into \[x=k/20\] to get the answer you were looking for!

OpenStudy (darkbluechocobo):

5=250/20

OpenStudy (darkbluechocobo):

5=125

OpenStudy (anonymous):

Do you mean x instead of 5?

OpenStudy (anonymous):

I think I see where you're getting confused. You have two pairs of numbers: 1. x=5, y=50 2. x=?, y=20 You can't mix up these pairs because the value of x depends on y. So you used to first pair to solve for k. Now use the second pair to solve for x. Does this help?

OpenStudy (darkbluechocobo):

ohhhhh

OpenStudy (darkbluechocobo):

x=250/50 x=12.5

OpenStudy (anonymous):

Nice! You're correct lol

OpenStudy (darkbluechocobo):

buht thats not an answer T-T

OpenStudy (anonymous):

xD what is the answer?

OpenStudy (anonymous):

Oh. You got the answer right. How would you write 12.5 as a fraction? Or think about how do you reduce the fraction \[ \frac{250}{20}\]

OpenStudy (darkbluechocobo):

25/2

OpenStudy (anonymous):

There you go!

OpenStudy (darkbluechocobo):

alright lol i just realized one more may i do it with you and see if i can get it without helps :p

OpenStudy (anonymous):

haha of course

OpenStudy (darkbluechocobo):

Find the missing value. y varies inversely as x. y = 10 when x = 5. Find y when x = 2.

OpenStudy (darkbluechocobo):

y=k/x 10=k/5 k=50

OpenStudy (anonymous):

Good

OpenStudy (darkbluechocobo):

y=50/2 y=25

OpenStudy (anonymous):

Yup. Easy!

OpenStudy (darkbluechocobo):

ohmai thank you I can dothem on my own nows :D

OpenStudy (darkbluechocobo):

10/10 heuheueheueheuehue

OpenStudy (anonymous):

haha you're very welcome :)